# Taipei 10617, Taiwan

### v2

@@ Abstract
The Lie algebra of symmetries generated by the left-moving current $j=\partial_{-}\phi$ in the $2d$ single scalar conformal field theory is infinite dimensional, exhibiting mutually commuting subalgebras. The infinite dimensional mutually commuting subalgebras define integrable deformations of the $2d$ single scalar conformal field theory which preserve the Poisson bracket structure. We study these mutually commuting subalgebras, finding general properties that the generators of such a subalgebra must satisfy. Along the way, we derive constraints on integrable equations of the Korteweg–de Vries type.
@@ 1 Introduction
The Korteweg–de Vries equation has a long and interesting history, first written down in 1871 by Boussinesq Boussinesq to approximate the behavior of solitary water waves after observations made by Russell russell1837experimental , and then further studied by Korteweg and de Vries Korteweg01051895 . The equation describes the elevation $h-u(x,t)$ ($-u$ small compared to $h$) of the surface of water in a rectangular channel at a horizontal position $x$ and time $t$. In a particular convention of units and parameters, the Korteweg–de Vries equation is
@@
Motivated to understand this peculiar property, the general exact solution to the Korteweg–de Vries equation was constructed via the inverse scattering transform Gardner:1967wc , and the equation was subsequently found to exhibit infinitely many commuting charges in a series of papers Kruskal1 ; Kruskal2 ; 10.1063/1.1664873 ; 10.1063/1.1665772 ; 10.1063/1.1665232 , including Laxorig where the concept of Lax pairs were introduced. Because of its infinitely many commuting charges, equations of this type are known as integrable equations, hearkening back to the notion of Liouville integrability of one dimensional Hamiltonian systems Liouville1855 .
Since then many integrable equations, or integrable field theories, have been found in $d=2$ spacetime dimensions, exhibiting particle-like “soliton” solutions. These solvable theories have offered a useful non-perturbative laboratory to understand phenomenologically relevant concepts e.g. PhysRevD.11.3424 ; PhysRevD.11.2088 ; ZAMOLODCHIKOV1979253 ; Dubovsky:2015zey ; Copetti:2024rqj . Beyond this, $2d$ integrable models appear commonly within string theory DHoker:1982wmk ; Metsaev:1998it ; PhysRevD.69.046002 ; Beisert_2011 , and more broadly in gauge theory Ablowitz1993 ; Costello_2018 ; Costello:2018gyb ; Costello:2019tri ; Gutperle:2014aja ; Beccaria:2015iwa . There are also deep connections between $2d$ integrable models and $2d$ conformal field theory Sasaki:1987mm ; Eguchi:1989hs ; ZAMOLODCHIKOV1989641 ; Bazhanov:1994ft ; Smirnov:2016lqw .
@@
$u_{t}-\left(V(\{u,u_{x},u_{xx},\dots\},x)\right)_{x}=0$ (1.3)
@@
[…]
where $\frac{\delta}{\delta u(x,t)}$ is a functional derivative with respect to $u(x,t)$. We are therefore searching for potentials $V$ which result in an integrable equation, having infinitely mutually commuting charges in the sense of (1.4).
Any potential $V=V(u)$ which does not depend on derivatives of $u$ or explicitly on $x$ results in an integrable equation, because any function $U(u)$ is an integrating factor, implying a conserved current for each such $U(u)$
@@
We are mainly concerned with potentials $V=V(\{u,u_{x},u_{xx},\dots\})$ that are allowed to depend on derivatives of $u$. In particular, we are concerned with potentials which are derivable from an action principle in a similar fashion to the Korteweg–de Vries equation, that are order $2n$ in $x$ derivatives of $u$. These potentials are always linear in $u_{2n}$, where $u_{i}$ is the $i$th $x$ derivative of $u$, which we write generically as
@@
where $A$ is an order $n$ functional of $u$, and $B$ is an order $12n-$ 1 functional of $u$, and is a polynomial in $u_{n+1},\dots,u_{2n-1}$. As an example, by defining $u=\phi_{x}$, the Korteweg–de Vries equation (1.1) can be obtained from the following action by varying with respect to $\phi$
[…]
@@
It was clear early on that not any Korteweg–de Vries type potential results in an integrable equation. For instance, it was observed that an equation with a potential of the form $V=\alpha u^{n}-u_{xx}$ is integrable only when $n=2,3$ Kruskal1 , with $n=3$ corresponding to the modified Korteweg–de Vries equation. By now there are several classifications of potentials of order 2 and 4 up to “almost invertible” transformations SISvinolupov_1992 using the so-called “symmetry approach” and the important concept of a formal recursion operator Svinolupov1982 ; Sokolov1985 ; SISvinolupov_1992 ; Mikhailov1991 ; Mikhailov2009 , summarized relatively recently in Heredero:2019arc with more references therein. A complete classification of all integrable equations of the type (1.3), or even the potentials (1.8), is still needed.
@@
The purpose of this paper is to further develop an alternative approach to classifying $2d$ integrable models introduced in Lindwasser:2024qyh , which can find within this approach necessary conditions on the lowest maximum nonzero order potential $V$ in each hierarchy, narrowing in on the parameter space of integrable equations of the type (1.8), and also other types of integrable equations. To do this, we reformulate the problem as a Lie algebra problem as in Lindwasser:2024qyh . As we explain in more detail in section 2, equations of the type (1.8) can be obtained from particular deformations of the $2d$ single scalar conformal field theory which preserve an infinite number of mutually commuting symmetries
@@
where ($\ell(\{j_{n}\},x^{-})$ is a functional of $j=\partial_{-}\phi$ and its $x^{-}$ derivatives up to order $n$. The equation of motion of this action takes the form (1.8) after making the identifications $x^{+}\to t$, $x^{-}\to x$, $j\to u$, and $\frac{\lambda}{2}\mathcal{E}(\ell)\to V$, where $\mathcal{E}(\ell)$ is the Euler–Lagrange equation of ($\ell(\{j_{n}\},x^{-})$, calculated as if it were a $1d$ Lagrangian of order $n$.
The Lie algebra of symmetries of the action (1.16) when $\lambda=0$ is infinite dimensional, with infinite dimensional mutually commuting subalgebras. We will see how each such subalgebra defines an integrable hierarchy, and so the problem at hand becomes finding all infinite dimensional mutually commuting subalgebras. In section 3, we will summarize the necessary conditions on these subalgebras found here and in Lindwasser:2024qyh .
The infinite dimensional mutually commuting subalgebras of the $2d$ single scalar conformal field theory generated by infinitesimal transformations $\delta_{i}$ such that $[\delta_{i},\delta_{j}]=0$ for all $i,j$ may admit integrable deformations, by which we mean deformations to the action $S\to S+S_{\text{deform}}$ satisfying $\delta_{i}S_{\text{deform}}=0$ for all $i$, which are not of the form (1.16), like the deformations defining the Liouville, sine-Gordon, and Bullough–Dodd (or Tzitzéica) models. In this way, studying these subalgebras is more general than classifying integrable equations of the type (1.8). One may think of this paper more simply as studying properties of the Lie algebra of symmetries present in the $2d$ single scalar conformal field theory. Constraints on integrable deformations are obtained from this analysis readily.
Formulating this problem through the Lie algebra of symmetries present in the $2d$ single scalar conformal field theory makes it clear that the same exercise can be performed on a much broader class of theories. In particular, any $2d$ action with a left(right)-moving current, like a collection of scalars, a collection of left(right)-moving fermions, Wess–Zumino–Witten models, or principal chiral models, has an infinite dimensional Lie algebra of symmetries. In the same way as the $2d$ single scalar conformal field theory, the infinite dimensional mutually commuting subalgebras of all of these theories define integrable hierarchies. By studying the Lie algebras associated with all such theories, one can find a large class of integrable models. We study here the case of the single scalar.
@@ 1.2 Outline
In section 2, we will review the Lie algebra of symmetries present in the classical $2d$ single scalar conformal field theory in terms of generators defined on the space of $1d$ Lagrangians ($\ell(\{j(x^{-}),\partial_{-}j(x^{-}),\dots\},x^{-})$, and show how its infinite dimensional mutually commuting subalgebras define integrable hierarchies of the type (1.8). Then in section 3, we summarize the necessary conditions on Korteweg–de Vries type hierarchies found here and in Lindwasser:2024qyh . In particular we describe the main result of this paper, which is the necessary $j_{n_{1}}$ dependence of the lowest nonzero order $n_{1}$ Lagrangian $\ell_{n_{1}}$ in any infinite dimensional mutually commuting subalgebra when $n_{1}=1,2,3,4$ and more generally when $n_{1}>4$ depending on certain technical details. In section 4, we provide a proof for this. Finally, in section 5, we discuss our results and present some questions for future research.
@@ 2 Preliminaries
The existence of a left-moving current $j(x^{-})$ is very constraining on the dynamics of the theory, because it implies the existence of infinitely many left-moving currents ($\ell(\{j(x^{-}),\partial_{-}j(x^{-}),\dots\},x^{-})$ that are arbitrary functionals of $j(x^{-})$. It is useful to assume that these functionals have a maximum finite order $k$ derivative of $j(x^{-})$ which the functional depends on. In this case, we will use the shorthand ($\ell(\{j_{k}(x^{-})\},x^{-})$ or more often simply as $\ell_{k}$, to indicate such a functional, where $j_{k}(x^{-})\equiv\partial_{-}^{k}j(x^{-})$. The Lie algebra is generated by the charges associated with these currents
[…]
Assuming that $j(x^{-})$ satisfies vanishing (at infinity) or periodic boundary conditions, unless it has a term that is independent of $j(x^{-})$, any functional which is a total derivative $\ell=\partial_{-}f(\{j_{k}(x^{-})\},x^{-})$ results in a vanishing charge. The set of functionals which generate the algebra then is the set of all functionals modulo total derivatives, or in other words the set of all $1d$ Lagrangians of $j(x^{-})$. The Lie algebra then acts on the vector space of $1d$ Lagrangians.
We will say that two $1d$ Lagrangians are equivalent $\ell\sim\ell^{\prime}$ if and only if they generate the same Euler–Lagrange equation, i.e. they differ by a total derivative $\ell=\ell^{\prime}+\partial_{-}f$. A $1d$ Lagrangian ($\ell(\{j_{k}(x^{-})\},x^{-})$ which depends on $j(x^{-})$ and its derivatives up to order $k$ may be equivalent to a Lagrangian ($\ell^{\prime}(\{j_{k^{\prime}}(x^{-})\},x^{-})$ with $k^{\prime}<k$. We define the order of a Lagrangian to be the minimum value of $k$ ($k$ being the maximum order derivative of $j(x^{-})$ that $\ell$ depends on) that can be obtained by adding total derivatives. We will assume without loss of generality that ($\ell(\{j_{k}(x^{-})\},x^{-})$ has been reduced such that $k$ represents the order of $\ell$.

### v3

@@ Abstract
The Lie algebra of symmetries generated by the left-moving current $j=\partial_{-}\phi$ in the $2d$ single scalar conformal field theory is infinite dimensional, exhibiting mutually commuting subalgebras. The infinite dimensional mutually commuting subalgebras define integrable deformations of the $2d$ single scalar conformal field theory which preserve the Poisson bracket structure. We study these mutually commuting subalgebras, finding general properties that the generators of such a subalgebra must satisfy. Along the way, we derive constraints on integrable equations of the Korteweg–de Vries type. We also confirm that the recently found [j]=0,1,2[j]=0,-1,-2 mutually commuting subalgebras are infinite dimensional.
@@ 1 Introduction
The Korteweg–de Vries equation has a long and interesting history, first written down in 1871 by Boussinesq Boussinesq to approximate the behavior of solitary water waves after observations made by Russell russell1837experimental , and then further studied by Korteweg and de Vries Korteweg01051895 . The equation describes the elevation hu(x,t)h-u(x,t) (u-u small compared to $h$) of the surface of water in a rectangular channel at a horizontal position $x$ and time $t$. In a particular convention of units and parameters, the Korteweg–de Vries equation is
@@
Motivated to understand this peculiar property, the general exact solution to the Korteweg–de Vries equation was constructed via the inverse scattering transform Gardner:1967wc , and the equation was subsequently found to exhibit infinitely many commuting charges in a series of papers Kruskal1 ; Kruskal2 ; 10.1063/1.1664873 ; 10.1063/1.1665772 ; 10.1063/1.1665232 , including Laxorig where the concept of Lax pairs were introduced. Because of its infinitely many commuting charges, equations of this type are known as integrable equations, harkening back to the notion of Liouville integrability of one dimensional Hamiltonian systems Liouville1855 .
Since then many integrable equations, or integrable field theories, have been found in $d=2$ spacetime dimensions, exhibiting particle-like “soliton” solutions. These solvable theories have offered a useful non-perturbative laboratory to understand phenomenologically relevant concepts, e.g., PhysRevD.11.3424 ; PhysRevD.11.2088 ; ZAMOLODCHIKOV1979253 ; Dubovsky:2015zey ; Copetti:2024rqj . Beyond this, $2d$ integrable models appear commonly within string theory DHoker:1982wmk ; Metsaev:1998it ; PhysRevD.69.046002 ; Beisert_2011 and, more broadly, in gauge theory Ablowitz1993 ; Costello_2018 ; Costello:2018gyb ; Costello:2019tri ; Gutperle:2014aja ; Beccaria:2015iwa . There are also deep connections between $2d$ integrable models and $2d$ conformal field theory Sasaki:1987mm ; Eguchi:1989hs ; ZAMOLODCHIKOV1989641 ; Bazhanov:1994ft ; Smirnov:2016lqw .
@@
$u_{t}-\left(V(\{u,u_{x},u_{xx},\dots\},x)\right)_{x}=0,$ $V=\frac{\delta H}{\delta u(x,t)}$ (1.3)
@@
$\{F,G\}\equiv\int dx\frac{\delta F}{\delta u(x,t)}\frac{d}{dx}\frac{\delta G}{\delta u(x,t)}$ (1.4)
where $\frac{\delta}{\delta u(x,t)}$ is a functional derivative with respect to $u(x,t)$. The functional $H$ in (1.3) is the Hamiltonian of this system with respect to this Poisson bracket. We are therefore searching for potentials $V$ that result in an integrable equation, having infinitely mutually commuting charges in the sense of (1.4).
Any potential $V=V(u)$ that does not depend on derivatives of $u$ or explicitly on $x$ results in an integrable equation, because any other function $U(u)$ is an integrating factor, implying a conserved current for each such $U(u)$,
@@
We are mainly concerned with potentials $V=V(\{u,u_{x},u_{xx},\dots\})$ that are allowed to depend on derivatives of $u$. In particular, we are concerned with potentials that are derivable from an action principle in a similar fashion to the Korteweg–de Vries equation, that are order $2n$ in $x$ derivatives of $u$. These potentials are always linear in $u_{2n}$, where $u_{i}$ is the $i$th $x$ derivative of $u$, which we write generically as
@@
where $A$ is an order $n$ functional of $u$, and $B$ is an order 2n12n-1 functional of $u$, and is a polynomial in $u_{n+1},\dots,u_{2n-1}$. As an example, by defining $u=\phi_{x}$, the Korteweg–de Vries equation (1.1) can be obtained from the following action by varying with respect to $\phi$
$S_{\text{KdV}}=\int dxdt\left(\phi_{t}\phi_{x}-\phi_{xx}^{2}-2\phi_{x}^{3}\right)$ (1.9)
@@
It was clear early on that not any Korteweg–de Vries type potential results in an integrable equation. For instance, it was observed that an equation with a potential of the form $V=\alpha u^{n}-u_{xx}$ is integrable only when $n=2,3$ Kruskal1 , with $n=3$ corresponding to the modified Korteweg–de Vries equation.
Demonstrating integrability for a generic Korteweg–de Vries type potential $V$ is generally difficult. The few examples so far mentioned exhibit nontrivial Lax pairs, whose existence prove integrability. Lax pairs, however, do not necessarily exist whenever an equation is integrable, and so other techniques for proving integrability are needed in this setting. Instead, the more general notion of a formal recursion operator $\Lambda$, which is a pseudo-differential operator that satisfies an equation $\Lambda_{t}=[V_{*},\Lambda]$, where $V_{*}$ is an operator associated with the potential, necessarily exists when $V$ results in an integrable equation. A psuedo-differential operator is an operator that is formally expanded as a sum of terms with derivatives up to some order $r$, including possible negative orders
$\Lambda=\sum_{n=-\infty}^{r}\Lambda_{n}(\{u,u_{x},u_{xx},\dots\},x)\partial_{x}^{n}$ (1.10)
As an example, the Korteweg–de Vries equation (1.1) has the formal recursion operator
$\Lambda_{\text{KdV}}=\partial_{x}^{2}-4u-2u_{x}\partial_{x}^{-1}$ (1.11)
whereas the modified Korteweg–de Vries equation, with a potential $V_{\text{mKdV}}=2u^{3}-u_{xx}$, has the formal recursion operator
$\Lambda_{\text{mKdV}}=\partial_{x}^{2}-4u^{2}-4u_{x}\partial_{x}^{-1}(u\,\cdot\,)$ (1.12)
Once one has a nontrivial formal recursion operator $\Lambda$, the infinitely many charges are obtained by taking (not necessarily integer) powers $\Lambda^{k}$ and extracting the coefficient in front of $\partial_{x}^{-1}$. For a generic Korteweg–de Vries type potential $V$, it is difficult to solve for $\Lambda$. An old approach known as the “symmetry approach” to classifying integrable equations of the type (1.3) uses the existence of $\Lambda$, solving explicitly for enough $\Lambda_{n}$, providing constraints on $V$ that reduce the list of possible integrable equations of a given order to a finite number up to “almost invertible” transformations SISvinolupov_1992 . By now there is a complete classification of potentials of order 2, and a classification of potentials of order 4 with constant separant [^1] A Korteweg–de Vries type potential of order $2n$ with constant separant takes the form $V(\{u,u_{x},\dots,u_{2n}\})=B(\{u,u_{x},\dots,u_{2n-1}\})-u_{2n}$. up to almost invertible transformations SISvinolupov_1992 ; Svinolupov1982 ; Sokolov1985 ; Mikhailov1991 ; Mikhailov2009 ; meshkov2013integrable , summarized relatively recently in Heredero:2019arc with more references therein. Integrability of a given potential $V$ can be proven after showing that the resulting equation is equivalent to one of the integrable equations on these lists. A complete classification of integrable equations of the type (1.3) is however still needed.
@@
The purpose of this paper is to further develop an alternative approach to classifying $2d$ integrable models introduced in Lindwasser:2024qyh , which can find within this approach necessary conditions on the lowest maximum nonzero order potential $V$ in each hierarchy, narrowing in on the parameter space of integrable equations of the type (1.8) and also other types of integrable equations. This approach is complimentary to the older “symmetry approach”, and we will see how we can arrive at new constraints on Korteweg–de Vries type potentials of any order without constant separant.
To do this, we reformulate the problem as a Lie algebra problem as in Lindwasser:2024qyh . As we explain in more detail in section 2, equations of the type (1.8) can be obtained from particular deformations of the $2d$ single scalar conformal field theory which preserve an infinite number of mutually commuting symmetries
@@
where $\ell(\{j_{n}\},x^{-})$ is a functional of $j=\partial_{-}\phi$ and its $x^{-}$ derivatives up to order $n$. The equation of motion of this action takes the form (1.8) after making the identifications $^{+}\to t, xx$$^{-}\to x$, $j\to u$, and $\frac{\lambda}{2}\mathcal{E}(\ell)\to V$, where $\mathcal{E}(\ell)$ is the Euler–Lagrange equation of $\ell(\{j_{n}\},x^{-})$, calculated as if it were a $1d$ Lagrangian of order $n$.
The Lie algebra of symmetries of the action (1.19) when $\lambda=0$ is infinite dimensional, with infinite dimensional mutually commuting subalgebras. We will see how each such subalgebra defines an integrable hierarchy, and so the problem at hand becomes finding all infinite dimensional mutually commuting subalgebras. In a similar fashion to the symmetry approach of classifying integrable equations Heredero:2019arc , in the absence of simple structures that prove infinite dimensionality like the existence of a Lax pair, we will only be able to find necessary conditions on infinite dimensional mutually commuting subalgebras. In section 3, we will summarize the necessary conditions on these subalgebras found here and in Lindwasser:2024qyh .
The infinite dimensional mutually commuting subalgebras of the $2d$ single scalar conformal field theory generated by infinitesimal transformations $\delta_{i}$ such that $\delta_{i},\delta_{j}]=0$ for all $i,j$ may admit integrable deformations, by which we mean deformations to the action $S\to S+S_{\text{deform}}$ satisfying $\delta_{i}S_{\text{deform}}=0$ for all $i$, which are not of the form (1.19), like the deformations defining the Liouville, sine-Gordon, and Bullough–Dodd (or Tzitzéica) models. In this way, studying these subalgebras is more general than classifying integrable equations of the type (1.8). One may think of this paper more simply as studying properties of the Lie algebra of symmetries present in the $2d$ single scalar conformal field theory. Constraints on integrable deformations are then obtained from this analysis.
Formulating this problem through the Lie algebra of symmetries present in the $2d$ single scalar conformal field theory makes it clear that the same exercise can be performed on a much broader class of theories. In particular, any $2d$ action with a left(right)-moving current, like a collection of scalars, a collection of left- (right-) moving fermions, Wess–Zumino–Witten models, or principal chiral models, has an infinite dimensional Lie algebra of symmetries. In the same way as the $2d$ single scalar conformal field theory, the infinite dimensional mutually commuting subalgebras of all of these theories define integrable hierarchies. By studying the Lie algebras associated with all such theories, one can find a large class of integrable models. We study here the case of the single scalar.
@@ 1.2 Outline
In section 2, we will review the Lie algebra of symmetries present in the classical $2d$ single scalar conformal field theory in terms of generators defined on the space of $1d$ Lagrangians $\ell(\{j(x^{-}),\partial_{-}j(x^{-}),\dots\},x^{-})$, and show how its infinite dimensional mutually commuting subalgebras define integrable hierarchies of the type (1.8). We will also confirm in subsection 2.3 that the mutually commuting subalgebras corresponding to [j]=0,1,2[j]=0,-1,-2 found in Lindwasser:2024qyh are infinite dimensional. Then in section 3, we summarize the necessary conditions on Korteweg–de Vries type hierarchies found here and in Lindwasser:2024qyh . In particular we describe the main result of this paper, which is the necessary $j_{n_{1}}$ dependence of the lowest nonzero order $n_{1}$ Lagrangian $\ell_{n_{1}}$ in any infinite dimensional mutually commuting subalgebra when $n_{1}=1,2,3,4$ and, more generally, when $n_{1}>4$ depending on certain technical details. In section 4, we provide a proof for this. Finally, in section 5, we discuss our results and present some questions for future research.
@@ 2 Preliminaries
The existence of a left-moving current $j(x^{-})$ is very constraining on the dynamics of the theory, because it implies the existence of infinitely many left-moving currents $\ell(\{j(x^{-}),\partial_{-}j(x^{-}),\dots\},x^{-})$ that are arbitrary functionals of $j(x^{-})$. It is useful to assume that these functionals have a maximum finite order $k$ derivative of $j(x^{-})$ which the functional depends on. In this case, we will use the shorthand $\ell(\{j_{k}(x^{-})\},x^{-})$ or more often simply as $\ell_{k}$, to indicate such a functional, where $j_{k}(x^{-})\equiv\partial_{-}^{k}j(x^{-})$. The Lie algebra is generated by the charges associated with these currents
$Q_{\ell}=\int dx^{-}\ell(\{j_{k}(x^{-})\},x^{-})$ (2.2)
Assuming that $j(x^{-})$ satisfies vanishing (at infinity) or periodic boundary conditions, unless it has a term that is independent of $j(x^{-})$, any functional that is a total derivative $\ell=\partial_{-}f(\{j_{k}(x^{-})\},x^{-})$ results in a vanishing charge. The set of functionals that generate the algebra then is the set of all functionals modulo total derivatives or, in other words, the set of all $1d$ Lagrangians of $j(x^{-})$. The Lie algebra then acts on the vector space of $1d$ Lagrangians.
We will say that two $1d$ Lagrangians are equivalent $\ell\sim\ell^{\prime}$ if and only if they generate the same Euler–Lagrange equation, i.e., they differ by a total derivative $\ell=\ell^{\prime}+\partial_{-}f$. A $1d$ Lagrangian $\ell(\{j_{k}(x^{-})\},x^{-})$ that depends on $j(x^{-})$ and its derivatives up to order $k$ may be equivalent to a Lagrangian $\ell^{\prime}(\{j_{k^{\prime}}(x^{-})\},x^{-})$ with $k^{\prime}<k$. We define the order of a Lagrangian to be the minimum value of $k$ ($k$ being the maximum order derivative of $j(x^{-})$ that $\ell$ depends on) that can be obtained by adding total derivatives. As a simple example, $\ell(\{j_{6}(x^{-})\},x^{-})=j(x^{-})j_{6}(x^{-})$ is order 3, because $jj_{6}\sim-j_{1}j_{5}\sim j_{2}j_{4}\sim-j_{3}j_{3}$. We will assume without loss of generality that $\ell(\{j_{k}(x^{-})\},x^{-})$ has been reduced such that $k$ represents the order of $\ell$.